OpenStudy (eunice_o):

solve for x. z=3x-7y

OpenStudy (eunice_o):

@RhondaSommer

OpenStudy (eunice_o):

do we change this to y=mx+b?

OpenStudy (mathstudent55):

You need to keep the term with x where it is and get rid of all other terms from that side. On the right side, you have 3x - 7y. How do you get rid of -7y from the right side?

RhondaSommer (rhondasommer):

Z? are you sure it is Z?

RhondaSommer (rhondasommer):

ohhh no I know what you mean

OpenStudy (eunice_o):

yes i am sure it is z

RhondaSommer (rhondasommer):

I would do what s/he said and add 7y to both sides

RhondaSommer (rhondasommer):

z+7y=3x-7y+7y

OpenStudy (eunice_o):

but we cant use the y=mx+b method ?

RhondaSommer (rhondasommer):

no you cant. unfortunately all you can do is get x alone

OpenStudy (eunice_o):

okay i did that but what i do with the 3x?

OpenStudy (eunice_o):

i divide all the number by 3?

RhondaSommer (rhondasommer):

yep!

OpenStudy (eunice_o):

then i cancel out the 3 then i get x=z/3 +7y/3?

RhondaSommer (rhondasommer):

\[z+7y=3x\]\[\frac{z+7y}{3}=\frac{3x}{3}\]\[\frac{z+7y}{3}=x\]

RhondaSommer (rhondasommer):

so yep :D

OpenStudy (eunice_o):

:)